# ZPressor: Bottleneck-Aware Compression for Scalable Feed-Forward 3DGS

### v2

@@ Abstract
Feed-forward 3D Gaussian Splatting (3DGS) models have recently emerged as a promising solution for novel view synthesis, enabling one-pass inference without the need for per-scene 3DGS optimization. However, their scalability is fundamentally constrained by the limited capacity of their encoders, leading to degraded performance or excessive memory consumption as the number of input views increases. In this work, we analyze feed-forward 3DGS frameworks through the lens of the Information Bottleneck principle and introduce ZPressor, a lightweight architecture-agnostic module that enables efficient compression of multi-view inputs into a compact latent state $Z$ that retains essential scene information while discarding redundancy. Concretely, ZPressor enables existing feed-forward 3DGS models to scale to over 100 input views at 480P resolution on an 80GB GPU, by partitioning the views into anchor and support sets and using cross attention to compress the information from the support views into anchor views, forming the compressed latent state $Z$. We show that integrating ZPressor into several state-of-the-art feed-forward 3DGS models consistently improves performance under moderate input views and enhances robustness under dense view settings on two large-scale benchmarks DL3DV-10K and RealEstate10K. The video results, code and trained models are available on our project page: https://lhmd.top/zpressor.
@@ 1 Introduction
To address this limitation, feed-forward 3DGS [7, 8] has been introduced to improve the usability of 3DGS. Unlike conventional 3DGS approaches that rely on slow per-scene backward optimization, feed-forward 3DGS introduces an “encoder” to extract scene dependent features from input images, allowing the model to benefit from large scale training and predict 3DGS in a single forward pass. Despite notable progress [9, 10, 11, 12, 13], these methods remain constrained to a small number of input views, limiting their ability to fully utilize datasets with dense multiple input views [14, 15, 16, 17]. For example, our experiments show that the state-of-the-art model DepthSplat [12] suffers a significant performance drop and increased computational cost as input views become denser (see Tab. 1 and Fig. 1), . While better engineering might alleviate this memory issue to some extent, it cannot address the huge performance degradation. Upon examining the architecture of several representative feed-forward 3DGS models [7, 8, 12], we identify the limited capacity of the encoder as the root cause. By design, it struggles to scale with denser input due to information overload and high computational cost.
Rather than introducing yet another ad-hoc encoder design, this work revisits the feed-forward 3DGS framework with inspiration from the Information Bottleneck (IB) [18] principle. IB offers a theoretical foundation for learning compact representations that preserve only task-relevant information. In the context of NVS with dense input views, we hypothesize that a latent representation can be learned to capture essential scene information while discarding redundant details in dense multi-view inputs. By encouraging the formation of such a compressed yet informative representation, we aim to improve the scalability of feed-forward 3DGS models (detailed in Sec. 3.2). Building on this insight, we propose a lightweight module, termed ZPressor, designed to be seamlessly integrated into the encoder of existing feed-forward 3DGS models to enhance their scalability. Unlike typical efforts that might rely on engineering optimizations such as memory-efficient attention [19] or activation checkpointing [20], our approach adopts a principled perspective grounded in representation learning, aiming to address core architectural limitations under dense input views settings.
@@
To validate the effectiveness of ZPressor, we integrate ZPressor into several state-of-the-art feed forward 3DGS models, including pixelSplat [7], MVSplat [8], and DepthSplat [12], and conduct extensive experiments on large-scale benchmarks such as DL3DV-10K [17] and RealEstate-10K [21]. Results show that integrating ZPressor consistently boosts the performance of baseline models under a moderate number of input views (e.g., 12 views), and helps them maintain reasonable accuracy and computational cost even with very dense inputs (e.g., 36 views, as shown in Fig. 1), where the original models typically degrade dramatically or run out of memory. Our contributions are threefold:
@@ 2 Related Work
Information bottleneck and its applications. The challenge of managing and processing vast quantities of information is a central theme in the development of large-scale machine learning models, particularly in the visual domain [22, 23, 24, 25]. The Information Bottleneck principle [18] formalizes the problem of extracting a compressed representation $Z$ from the input $X$, such that $Z$ is maximally informative about the target $Y$. The IB principle was subsequently extended to the domain of deep learning [26, 27, 28], the Deep Variational Information Bottleneck [29] providing a tractable lower bound on the IB objective, bridging the gap between the theoretical IB principle and practical deep learning applications. A series of works have applied the IB principle to multi-view inputs [30, 31, 32], extract information that is common or shared across multiple views while discarding view-specific or redundant information. The drive towards efficient 3D scene reconstruction, especially within the context of 3DGS [1], has also seen the adoption of information-theoretic ideas. StreamGS [33] tackle redundancy in image streams by merging superfluous Gaussians through cross-frame feature aggregation, while other works [34, 35] focus on compressing the learned 3D Gaussians. While demonstrating the effectiveness and necessity of compression in multi-view data and 3D reconstruction, none of the existing works have explored it in the context of feed-forward 3DGS. Our work aims to bridge this gap by introducing the lens of IB for information compression in this area.
@@
Optimization-based NeRF and 3DGS. Traditional novel view synthesis (NVS) methods primarily rely on image blending techniques [36, 37]. More recently, neural network-based approaches [38, 39, 40, 41] have advanced NVS by integrating it with deep learning models. In particular, NeRF [41] employs an MLP to map 3D spatial locations and viewing directions to radiance color and volume density. Numerous works [42, 43, 44, 45, 46, 47] have sought to improve NeRF’s efficiency and reconstruction quality. However, its reliance on volume rendering [48] hinders rendering speed, limiting its practicality in real-world applications. Recently, 3D Gaussian Splatting (3DGS)[1] and its variants[2, 49, 50, 51, 35] have emerged as efficient solutions for large-scale scene reconstruction and synthesis, offering explicit representations and fast rasterization-based rendering that outperform NeRF’s slower volumetric approach. Nonetheless, its requirement for slow per-scene optimization still poses challenges for deployment in downstream tasks.
@@
Feed-Forward NeRF and 3DGS. To address the limitation of slow per-scene optimization, PixelNeRF [52] pioneered the feed-forward NeRF (a.k.a. generalizable NeRF) by introducing an additional network that directly encodes input views into a NeRF representation. This allows the model to benefit from large-scale training and predict a scene representation in a single forward pass. This direction has since seen significant advancements [53, 54, 55, 56, 57], offering a promising path toward practical NeRF deployment. This paradigm has recently been extended to real-time rendering with 3DGS [1] replacing NeRF [41]. Among them, pixelSplat [7] pioneered the feed-forward 3DGS approach by combining epipolar transformers with depth prediction to predict 3D Gaussians from two input views. MVSplat [8] proposed an efficient cost-volume-based fusion strategy to improve multi-view reconstruction, while DepthSplat [12] leveraged monocular depth features to better recover fine 3D structures from sparse inputs. Despite the growing number of feed-forward 3DGS models [58, 59, 60, 61, 62], most follow a pixel-aligned design, where the number of 3D Gaussians scales linearly with the number of input views. This leads to significant memory and computational overhead as input views increase. While works like FreeSplat [10] and GGN [11] attempt to reduce the number of Gaussians by merging them via cross-view projection checking, they lack a principled framework. In contrast, our work provides a theoretical perspective on the information overload problem in feed-forward 3DGS by introducing the IB [18] principle. And we propose an architecture-agnostic module ZPressor that can be seamlessly integrated into existing feed-forward 3DGS models to improve performance under dense input view settings.
@@ 3.1 Overview of ZPressor
Our ZPressor is a plug-and-play module for compressing the multi-view inputs of feed-forward 3DGS, as illustrated in Fig. 2. Formally, given $K$ input views $\mathcal{V}=\{{\mathbf{V}}_{i}\}_{i=1}^{K}$ where ${\mathbf{V}}_{i}\in\mathbb{R}^{H\times W\times 3}$ and their corresponding camera poses $\mathcal{P}=\{{\mathbf{P}}_{i}\}_{i=1}^{K}$, our ZPressor takes the extracted features from each view as input:
@@
Figure 2: Overview of ZPressor for Feed-Forward 3DGS. Our proposed ZPressor is a plug-and-play module designed for feed-forward 3DGS frameworks. It addresses the challenge of processing dense input views by strategically grouping input view features ${\mathcal{X}}$ based on selected anchor views, then features within each respective group are compressed as ${\mathcal{Z}}$.
@@ 3.2 Information Analysis of Feed-Forward 3DGS
Existing feed-forward 3D Gaussian Splatting networks suffer from a dramatic performance drop and an exponential increase in computational cost when the amount of input view information grows (see  Tab. 1), primarily due to information redundancy and the lack of adaptive information compression mechanisms. Specifically, the total information of the scene is represented by the joint entropy $H(\mathbf{F}_{1},\mathbf{F}_{2},...,\mathbf{F}_{K})$, which is not merely the sum of the individual entropies of the features from all views, i.e., $\sum_{i=1}^{K}H(\mathbf{F}_{i})$. Therefore, there is a significant amount of redundant information in the features, and it is crucial to remove the irrelevant information after feature extraction while preserving its predictive power, which allows for the efficient use of information from the input views.
A priciple way for modeling this is the Information Bottleneck (IB) [18], which minimizes the IB scores as
@@
As shown in Fig. 2, the Prediction Score is typically modelled by the Gaussian predictor. In the next section, we introduce ZPressor, a novel module applied concatenated with the image encoder to model the Compression Score through three consecutive designs, yielding a compact presentation.
@@ 3.3 Information Compression Module: ZPressor
Anchor view selection (Fig. 2 Step 1). Given a set of camera positions $\mathcal{T}=\{\mathbf{T}_{i}\}_{i=1}^{K}\in\mathbb{R}^{K\times 3}$ calculated from camera parameters $\mathcal{P}$, where $f\colon\mathbf{T}\to\mathbf{F}$ such that $f$ is a bijective mapping, we first add a random view to the anchor view list $\mathcal{S}=\{\mathbf{T}_{a_{1}}\}$, where $\mathbf{T}_{a_{1}}\sim\text{Uniform}(\mathcal{T})$.
@@
Support-to-anchor assignment (Fig. 2 Step 2). Once anchors are selected, each support view is assigned to its nearest anchor based on camera position. This ensures that support views, which capture complementary scene details, are grouped with the most spatially relevant anchor views, thereby ensuring the effectiveness of information fusion. Formally, the cluster assignment to the $i$-th anchor view can be denoted as:
@@
Views information fusion (Fig. 2 Step 3). Upon obtaining the anchor-support view-based clusters, we aim to fuse the information within each cluster. The fusion mechanism should satisfy the following properties: 1) use the anchor views as the base, while effectively integrating information from the support views to enhance them, and 2) capture the similarity between the two sets of views, maintaining compactness while avoiding redundancy.
@@
${\mathcal{Z}}=\text{Cross-Attn}(Q,K,V),\quad Q\leftarrow\mathcal{X}_{\mathrm{anchor}},\quad K,V\leftarrow\mathcal{X}_{\mathrm{support}},$ (6)
@@
Training. Upon obtaining the compressed information ${\mathcal{Z}}$, we can apply the IB principle to regularize the information flow. Recall that, as in Eq. 3, our current goal should be to minimize the Compression Score $I({\mathcal{X}_{\mathrm{anchor}},\mathcal{X}_{\mathrm{support}}};{\mathcal{Z}})$ and maximize Prediction Score $I({\mathcal{Z}};{\mathcal{Y}})$. For the Compression Score, we apply a constraint on its complexity, i.e., setting the number of anchor views $N$ to a value acceptable for training. For the Prediction Score, we have
@@
where $H({\mathcal{Y}})$ is a constant representing the 3D Gaussians that model the underlying 3D scene. Hence, maximizing $I({\mathcal{Z}};{\mathcal{Y}})$ is equivalent to minimizing $H({\mathcal{Y}}\mid{\mathcal{Z}})$, which essentially encourages the predicted 3D Gaussians ${\mathcal{Y}}$ to resemble the original scene. Then we incorporate the IB principle into feed-forward 3DGS training, using the efficient variational estimate of Eq. 3 following [29, 26]. Since 3DGS is typically trained by rendering and supervising on the 2D image space, we optimize the feed-forward 3DGS with ZPressor by the variational rendering loss, e.g., MSE and LPIPS.
In our implementation, we append an additional self-attention layer to further enhance information flow within each cluster. In addition, we stack several blocks (each containing both cross- and self-attention) to further improve the effectiveness of the IB principle. These two general-purpose engineering techniques help boost performance, as shown in Tab. 3. A comprehensive description of ZPressor’s architecture is available in the supplementary material.

### v4

@@ Abstract
Feed-forward 3D Gaussian Splatting (3DGS) models have recently emerged as a promising solution for novel view synthesis, enabling one-pass inference without the need for per-scene 3DGS optimization. However, their scalability is fundamentally constrained by the limited capacity of their models, leading to degraded performance or excessive memory consumption as the number of input views increases. In this work, we analyze feed-forward 3DGS frameworks through the lens of the Information Bottleneck principle and introduce ZPressor, a lightweight architecture-agnostic module that enables efficient compression of multi-view inputs into a compact latent state $Z$ that retains essential scene information while discarding redundancy. Concretely, ZPressor enables existing feed-forward 3DGS models to scale to over 100 input views at 480P resolution on an 80GB GPU, by partitioning the views into anchor and support sets and using cross attention to compress the information from the support views into anchor views, forming the compressed latent state $Z$. We show that integrating ZPressor into several state-of-the-art feed-forward 3DGS models consistently improves performance under moderate input views and enhances robustness under dense view settings on two large-scale benchmarks DL3DV-10K and RealEstate10K. The video results, code and trained models are available on our project page: https://lhmd.top/zpressor.
@@ 1 Introduction
To address this limitation, feed-forward 3DGS [7, 8] has been introduced to improve the usability of 3DGS. Unlike conventional 3DGS approaches that rely on slow per-scene backward optimization, feed-forward 3DGS introduces an “encoder” to extract scene dependent features from input images, allowing the model to benefit from large scale training and predict 3DGS in a single forward pass. Despite notable progress [9, 10, 11, 12, 13], these methods remain constrained to a small number of input views, limiting their ability to fully utilize datasets with dense multiple input views [14, 15, 16, 17]. For example, our experiments show that the state-of-the-art model DepthSplat [12] suffers a significant performance drop and increased computational cost as input views become denser (see Tab.˜1 and Fig. 1), . While better engineering might alleviate this memory issue to some extent, it cannot address the huge performance degradation. Upon examining the architecture of several representative feed-forward 3DGS models [7, 8, 12], we identify the limited capacity of the whole network from image to 3DGS as the root cause.By design, it struggles to scale to denser inputs due to representation overload from excessive feature tokens and the resulting high computational cost.
Rather than introducing yet another ad-hoc encoder design, this work revisits the feed-forward 3DGS framework with inspiration from the Information Bottleneck (IB) [18] principle. IB offers a theoretical foundation for learning compact representations that preserve only task-relevant information. In the context of NVS with dense input views, we hypothesize that a latent representation can be learned to capture essential scene information while discarding redundant details in dense multi-view inputs. By encouraging the formation of such a compressed yet informative representation, we aim to improve the scalability of feed-forward 3DGS models (detailed in Sec.˜3.2). Building on this insight, we propose a lightweight module, termed ZPressor, designed to be seamlessly integrated into the encoder of existing feed-forward 3DGS models to enhance their scalability. Unlike typical efforts that might rely on engineering optimizations such as memory-efficient attention [19] or activation checkpointing [20], our approach adopts a principled perspective grounded in representation learning, aiming to address core architectural limitations under dense input views settings.
@@
To validate the effectiveness of ZPressor, we integrate ZPressor into several state-of-the-art feed forward 3DGS models, including pixelSplat [7], MVSplat [8], and DepthSplat [12], and conduct extensive experiments on large-scale benchmarks such as DL3DV-10K [17] and RealEstate-10K [21]. Results show that integrating ZPressor consistently boosts the performance of baseline models under a moderate number of input views (e.g., 12 views), and helps them maintain reasonable accuracy and computational cost even with very dense inputs (e.g., 36 views, as shown in Fig.˜1), where the original models typically degrade dramatically or run out of memory. Our contributions are threefold:
@@ 2.1 Information bottleneck and its applications
The challenge of managing and processing vast quantities of information is a central theme in the development of large-scale machine learning models, particularly in the visual domain [22, 23, 24, 25]. The Information Bottleneck principle [18] formalizes the problem of extracting a compressed representation $Z$ from the input $X$, such that $Z$ is maximally informative about the target $Y$. The IB principle was subsequently extended to the domain of deep learning [26, 27, 28, 29], VIB [30] providing a tractable lower bound on the IB objective, bridging the gap between the theoretical IB principle and practical deep learning applications. A series of works have applied the IB principle to multi-view inputs [31, 32, 33], extract information that is common or shared across multiple views while discarding view-specific or redundant information. The drive towards efficient 3D scene reconstruction, especially within the context of 3DGS [1], has also seen the adoption of information-theoretic ideas. StreamGS [34] tackle redundancy in image streams by merging superfluous Gaussians through cross-frame feature aggregation, while other works [35, 36] focus on compressing the learned 3D Gaussians. While demonstrating the effectiveness and necessity of compression in multi-view data and 3D reconstruction, none of the existing works have explored it in the context of feed-forward 3DGS. Our work aims to bridge this gap by introducing the lens of IB for information compression in this area.
2.2 Optimization-based NeRF and 3DGS
@@
Traditional novel view synthesis (NVS) methods primarily rely on image blending techniques [37, 38]. More recently, neural network-based approaches [39, 40, 41, 42] have advanced NVS by integrating it with deep learning models. In particular, NeRF [42] employs an MLP to map 3D spatial locations and viewing directions to radiance color and volume density. Numerous works [43, 44, 45, 46, 47, 48] have sought to improve NeRF’s efficiency and reconstruction quality. However, its reliance on volume rendering [49] hinders rendering speed, limiting its practicality in real-world applications. Recently, 3D Gaussian Splatting (3DGS)[1] and its variants[2, 50, 51, 52, 36] have emerged as efficient solutions for large-scale scene reconstruction and synthesis, offering explicit representations and fast rasterization-based rendering that outperform NeRF’s slower volumetric approach. Nonetheless, its requirement for slow per-scene optimization still poses challenges for deployment in downstream tasks.
2.3 Feed-Forward NeRF and 3DGS
@@
To address the limitation of slow per-scene optimization, PixelNeRF [53] pioneered the feed-forward NeRF (a.k.a. generalizable NeRF) by introducing an additional network that directly encodes input views into a NeRF representation. This allows the model to benefit from large-scale training and predict a scene representation in a single forward pass. This direction has since seen significant advancements [54, 55, 56, 57, 58], offering a promising path toward practical NeRF deployment. This paradigm has recently been extended to real-time rendering with 3DGS [1] replacing NeRF [42]. Among them, pixelSplat [7] pioneered the feed-forward 3DGS approach by combining epipolar transformers with depth prediction to predict 3D Gaussians from two input views. MVSplat [8] proposed an efficient cost-volume-based fusion strategy to improve multi-view reconstruction, while DepthSplat [12] leveraged monocular depth features to better recover fine 3D structures from sparse inputs. Despite the growing number of feed-forward 3DGS models [59, 60, 61, 62, 63], most follow a pixel-aligned design, where the number of 3D Gaussians scales linearly with the number of input views. This leads to significant memory and computational overhead as input views increase. While works like FreeSplat [10] and GGN [11] attempt to reduce the number of Gaussians by merging them via cross-view projection checking, they lack a principled framework. In contrast, our work provides a theoretical perspective on the representation overload problem in feed-forward 3DGS by introducing the IB [18] principle. And we propose an architecture-agnostic module ZPressor that can be seamlessly integrated into existing feed-forward 3DGS models to improve performance under dense input view settings.
@@ 3.1 Overview of ZPressor
Our ZPressor is an architecture-agnostic module for compressing the multi-view inputs of feed-forward 3DGS, as illustrated in Fig.˜2. Formally, given $K$ input views $\mathcal{V}=\{{\mathbf{V}}_{i}\}_{i=1}^{K}$ where ${\mathbf{V}}_{i}\in\mathbb{R}^{H\times W\times 3}$ and their corresponding camera poses $\mathcal{P}=\{{\mathbf{P}}_{i}\}_{i=1}^{K}$, our ZPressor takes the extracted features from each view as input:
@@
Figure 2: Overview of ZPressor for Feed-Forward 3DGS. Our proposed ZPressor is an architecture-agnostic module designed for feed-forward 3DGS frameworks. It addresses the challenge of processing dense input views by strategically grouping input view features $\mathcal{X}$} based on selected anchor views, then features within each respective group are compressed as $\mathcal{Z}$}.
@@ 3.2 Information Analysis of Feed-Forward 3DGS
Existing feed-forward 3D Gaussian Splatting networks suffer from a dramatic performance drop and an exponential increase in computational cost when the amount of input view information grows (see  Tab.˜1), primarily due to information redundancy and the lack of adaptive information compression mechanisms. Specifically, the total information of the scene is represented by the joint entropy $H(\mathbf{F}_{1},\mathbf{F}_{2},...,\mathbf{F}_{K})$, which is not merely the sum of the individual entropies of the features from all views, i.e., $\sum_{i=1}^{K}H(\mathbf{F}_{i})$. Therefore, there is a significant amount of redundant information in the features, and it is crucial to remove the irrelevant information after feature extraction while preserving its predictive power, which allows for the efficient use of information from the input views.
A priciple way for modeling this is the Information Bottleneck (IB) [18], which minimizes the IB scores as:
@@
As shown in Fig.˜2, the Prediction Score is typically modelled by the Gaussian predictor. In the next section, we introduce ZPressor, a novel module applied concatenated with the image encoder to model the Compression Score through three consecutive designs, yielding a compact presentation.
@@ 3.3 Information Compression Module: ZPressor
Anchor view selection (Fig.˜2 Step 1). Given a set of camera positions $\mathcal{T}=\{\mathbf{T}_{i}\}_{i=1}^{K}\in\mathbb{R}^{K\times 3}$ calculated from camera parameters $\mathcal{P}$, where $f\colon\mathbf{T}\to\mathbf{F}$ such that $f$ is a bijective mapping, we first add a random view to the anchor view list $\mathcal{S}=\{\mathbf{T}_{a_{1}}\}$, where $\mathbf{T}_{a_{1}}\sim\text{Uniform}(\mathcal{T})$.
@@
Support-to-anchor assignment (Fig.˜2 Step 2). Once anchors are selected, each support view is assigned to its nearest anchor based on camera position. This ensures that support views, which capture complementary scene details, are grouped with the most spatially relevant anchor views, thereby ensuring the effectiveness of information fusion. Formally, the cluster assignment to the $i$-th anchor view can be denoted as:
@@
Views information fusion (Fig.˜2 Step 3). Upon obtaining the anchor-support view-based clusters, we aim to fuse the information within each cluster. The fusion mechanism should satisfy the following properties: 1) use the anchor views as the base, while effectively integrating information from the support views to enhance them, and 2) capture the similarity between the two sets of views, maintaining compactness while avoiding redundancy.
@@
${\mathcal{Z}}=\text{Attention}(Q,K,V),\quad Q\leftarrow\mathcal{X}_{\mathrm{anchor}},\quad K,V\leftarrow\mathcal{X}_{\mathrm{support}},$ (6)
@@
Training. Upon obtaining the compressed information $\mathcal{Z}$}, we can apply the IB principle to regularize the information flow. Recall that, as in Eq.˜3, our current goal should be to minimize the Compression Score $I({\mathcal{X}_{\mathrm{anchor}},\mathcal{X}_{\mathrm{support}}};{\mathcal{Z}})$ and maximize Prediction Score $I({\mathcal{Z}};{\mathcal{Y}})$. For the Compression Score, we apply a constraint on its complexity, i.e., setting the number of anchor views $N$ to a value acceptable for training. For the Prediction Score, we have
@@
where $H({\mathcal{Y}})$ is a constant representing the 3D Gaussians that model the underlying 3D scene. Hence, maximizing $I({\mathcal{Z}};{\mathcal{Y}})$ is equivalent to minimizing $H({\mathcal{Y}}\mid{\mathcal{Z}})$, which essentially encourages the predicted 3D Gaussians $\mathcal{Y}$} to resemble the original scene.
Then we incorporate the IB principle into feed-forward 3DGS training, using the efficient estimate of Eq.˜3 following [30, 26]. Formally, the training objective is defined as:
$\mathcal{L}=\underset{\mathcal{Z}\sim p_{\theta}(\mathcal{Z}\mid\mathcal{X})}{\mathbb{E}}\bigl[-\log q_{\phi}(\mathcal{Y}\mid\mathcal{Z})\bigr]+\beta\,\underset{\mathcal{X}}{\mathbb{E}}\bigl[\text{KL}\bigl[p_{\theta}(\mathcal{Z}\mid\mathcal{X}),\,r(\mathcal{Z})\bigr]\bigr]$ (8)
where $\phi$ denotes the parameters of 3D Gaussians prediction network and $\log q_{\phi}(\mathcal{Y}\mid\mathcal{Z})$ can be modeled by the rendering loss (such as the MSE and LPIPS loss), $\theta$ denotes the parameters of the network preceding $\mathcal{Z}$ and $p_{\theta}(\mathcal{Z}\mid\mathcal{X})$ is the posterior probability estimate of $\mathcal{Z}$, $r(\mathcal{Z})\sim\mathcal{N}(\mathcal{Z}\mid\mu_{0},\Sigma_{0})$ is the Gaussian prior of $\mathcal{Z}$.
In our implementation, we append an additional self-attention layer to further enhance information flow within each cluster. In addition, we stack several blocks (each containing both cross- and self-attention) to further improve the effectiveness of the IB principle. These two general-purpose engineering techniques help boost performance, as shown in Tab.˜3.